OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (anonymous):

-3x-9=-3x-3-5 -3x-9=-3x-8 -3x=-3x+1 x=1/3 So maybe all real number

OpenStudy (anonymous):

@Cryxz that is incorrect because if you add 3x to both sides, the equation becomes 0 = 1, which is obvioously incorrect so the answer would be No Solution.

OpenStudy (anonymous):

think you guys for the help :)